Fermi National Laboratory
 

Accelerator Update for October 18, 2006 - October 20, 2006

 

Wednesday October 18

The day shift began with the Tevatron (TeV) and Antiproton Source (Pbar) off due to magnet replacements, with the Recycler, Linac, Booster, and the Main Injector (MI) operating normally, with MiniBooNE and NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) in access.

Technicians reported at 9:40 AM that the Booster RF Cavity was installed and ready for alignment.

Operations prepared Pbar for an access at 11:24 AM.

Safety system techs completed their test of MI at 11:58 AM.

Cryo system techs began cooling TeV sector A4 at 12:48 PM; sector B1 cooldown will follow.

Booster came back on line at 4:06 PM.

MI came back on line at 4:25 PM.

Operations established beam to MiniBooNE at 4:58 PM.

Operations established beam to the Pbar Debuncher at 5:59 PM. Experts then conducted two hours of studies.

Recycler experts stashed protons for studies at 6:41 PM.

Operations established beam to NuMI at 8:39 PM.

Thursday October 19

The midnight shift began with the TeV in access, and with NuMI and MiniBooNE taking beam.

At 12:43 AM, the NuMI horn tripped off for the first of three times in an hour on a ground fault indication. It reset each time.

Pbar studies ended at 1:39 AM.

Operations prepared Pbar for access at 8:25 AM.

Operations reported at 11:51 AM that the Neutron Therapy Facility (NTF) had completed treating patients for the day.

Operations verified that Pbar could stack antiprotons at 1:12 PM. Pbar experts then took reverse protons for studies.

Recycler experts completed their studies at 4:20 PM.

Safety system techs completed their test of CDF at 4:57 PM.

Safety system techs completed their test of D0 at 5:41 PM.

At 8:19 PM, Operations reported that a Debuncher power supply (D:QD) tripped off for the first of two times in twenty minutes on a ground fault. This did not affect the studies. Operations contacted an expert.

Friday October 20

The midnight shift began with the TeV in access, and with NuMI and MiniBooNE taking beam.

At midnight, a cold compressor for the TeV sector B3 area tripped off. Operators investigated, but left the compressor off; it will be checked in the morning.

The Pbar studies ended at 1:38 AM.

At 2:13 AM, the NuMI horn tripped off on a ground fault indication. It reset okay.

Operations prepared Pbar for access at 6:26 AM.

At 8:07 AM, a power supply expert reported that a Pbar magnet (D6Q18) might have a ground fault. The investigation continued.

The Plans for Friday and the Weekend

The TeV is on schedule to begin the D sector cooldown on Saturday night. Startup might be delayed by the results of the Pbar D6Q18 investigations.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/20/2006    email Fermilab